OpenStudy (anonymous):

Graph the equation with a diameter that has endpoints at (-2, -2) and (4, 2). Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Heey, this is kinda the same question as befre, but I don't think my answer for distance is correct

OpenStudy (anonymous):

@Directrix

OpenStudy (jdoe0001):

hmm what did you get for the distance?

OpenStudy (anonymous):

About 7.21

OpenStudy (jdoe0001):

hmm is there a radical form for it?

OpenStudy (anonymous):

Sorry...?

OpenStudy (jdoe0001):

well. you used the distance formula to get it.... so... that means you'd have a root of it something like \(\bf distance = \sqrt{\qquad}\)

OpenStudy (anonymous):

\[Distance= \sqrt{(-2-4)^2+(-2-2)^2}\]

OpenStudy (jdoe0001):

yeap... I got the same, which in decimal form is about 7.21 so, that's correct \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -2}}\quad ,&{\color{blue}{ -2}})\quad % (c,d) &({\color{red}{ 4}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad % distance value d = \sqrt{6^2+4^2}\implies d=\sqrt{52}\impliedby diameter \\ \quad \\ radius=\cfrac{\sqrt{52}}{2}\)

OpenStudy (anonymous):

So the radius would be 3.60

OpenStudy (jdoe0001):

using the midpoint formula, the center is at 1,0 thus \(\large (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ 1}},{\color{blue}{ 0}})\qquad radius={\color{purple}{ \frac{\sqrt{52}}{2}}}\)

OpenStudy (anonymous):

Wouldn't the center be (1,1)

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

lemme recheck that

OpenStudy (anonymous):

Oops, nevermind, my bad, it is 0

OpenStudy (jdoe0001):

yeah.. is 1,0

OpenStudy (jdoe0001):

shoot.. a bit truncated...lemme redo that

OpenStudy (jdoe0001):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ 1}},{\color{blue}{ 0}})\qquad radius={\color{purple}{ \frac{\sqrt{52}}{2}}} \\ \quad \\ \quad \\ (x-{\color{brown}{ 1}})^2+(y-{\color{blue}{ 0}})^2=\left( {\color{purple}{ \frac{\sqrt{52}}{2}}}\right)^2 \\ \quad \\ (x-1)^2+y=\cfrac{(\sqrt{52})^2}{2^2}\implies (x-1)^2+y=\cfrac{\cancel{52}}{\cancel{4}} \\ \quad \\ \implies (x-1)^2+y=13\)

OpenStudy (anonymous):

So the square root is cancelled off in the equation but I still have a radius of 3.60 right?

OpenStudy (jdoe0001):

if you want the decimal form of it, yes the advantage of the radical form is that, it doesn't have any missing floating values

OpenStudy (anonymous):

Yeah, but I need the decimal so that I can graph it

OpenStudy (jdoe0001):

but yes, 7.21 is a rounded up version, you lose a few decimals there half that is 3.6, also rounded up, again losing a few decimals

OpenStudy (jdoe0001):

welll. yes sure.. but yes, is 3.6 :)

OpenStudy (anonymous):

Thank u!!

OpenStudy (triciaal):